Fermi National Laboratory
 

Accelerator Update for April 13, 2007 - April 16, 2007

 

Friday April 13

The day shift began with the Operations monitoring Tevatron (TeV) store 5351 (68.04) and the Recycler stash (160.6E10), with the Antiproton Source (Pbar) stacking (26.1E10), with the Booster, Linac, Main Injector (MI), and Recycler operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 8:30 AM, the MiniBooNE autotune program got lost. Operators tuned the beam the old fashioned way and then handed it back to the program.

At 8:48 AM, a Linac RF station (LRF3) tripped off. It reset.

A MiniBooNE expert reported that a BPM module had failed and that was why the autotune program wasn’t working. Operations turned off MiniBooNE beam while experts investigated.

At 9:36 AM, Operations reported that the I- Source vacuum gauge had tripped off three times since 8 AM. It reset.

At 9:57 AM, the NuMI beam permit tripped off due to an exhaust fan tripping off. The fan came back on within five minutes and NuMI resumed taking beam.

At 10:03 AM, LRF1 tripped off for the first of four times in about three hours. There was no indication of what was causing the trips. It ran itself back up after each trip.

A MI expert began a two-hour study at 10:53 AM.

At 11:08 AM, MI tripped off for the first of four times in 20 minutes due to bad beam positions. It reset.

At 12:07 PM, the I- Source vacuum gauge tripped off for the first of two times in 30 minutes. It reset.

A MI expert began a slipstacking study at 1:09 PM.

At 2:20 PM, BRF6 tripped off. It reset.

At 2:37 PM, the current for the I- Source’s extractor magnet tripped off. This halted all beam. Experts investigated and replaced the magnet’s power supply. Beam resumed at 3:10 PM.

The MI slipstacking study ended at 3:35 PM.

At 3:57 PM, a Linac Klystron RF station (KRF3) tripped off. It reset.

Operations reported at 4 PM that the Recycler had stashed antiprotons twice during the day shift.

Operations terminated store 5351 at 5:21 PM.

Operations put the TeV into shot setup at 5:38 PM.

Operations established store 5352 at 7:32 PM with an initial luminosity of 198.5E30.

A TeV expert began a Tevatron Electron Lens (TEL1) study at 7:51 PM.

At 8:47 PM, a Booster vacuum pump (B:IPRF24) tripped off and caused BRF15 & 16 to trip off. Everything reset.

At 9:24 PM, a Booster vacuum pump (B:IPRF17) tripped off and caused BRF7 & 8 to trip off. Everything reset.

At 10:25 PM, a Meson MS2 area LCW water skid tripped off. Operators investigated, but could not turn the skid on. Experts will investigate further in the morning.

Operations reported at midnight that the Recycler had stashed three times during the evening shift.

Saturday April 14

The midnight shift began with Operations monitoring store 5352 (107.96E30) and the stash (108.6E10), with Pbar stacking (29.5E10), and with NuMI and MiniBooNE taking beam.

At 2:34 AM, LRF3 tripped off. It reset.

At 4:07 AM, LRF3 tripped off for the first of three times in about an hour. It reset.

At 5:43 AM, LRF5 tripped off. It reset.

Operations reported at 7:59 AM that the Recycler had stashed three times during the midnight shift.

At 8:34 AM, the I- Source vacuum gauge tripped off for the first of five times in three hours.

At 10:23 AM, BRF6 tripped off. It reset.

At 10:43 AM, Operations reported that the temperature for the Booster low-level RF room had alarmed high. The AC for the room was down and a portable AC unit had been installed, but it was breakered off. It was turned on.

At 11 AM, LRF5 tripped off. It reset.

At 12:32 PM, BRF6 tripped off for the first of three times in five and a half hours. It reset.

The Duty Mechanic reported at 2:04 PM that the MS2 LCW problem was due to a bad air compressor. They turned on the backup compressor and the water skid resumed normal operations.

An MI expert began a barrier bucket study at 2:50 PM. The study lasted three hours.

At 2:59 PM, MiniBooNE had another autotune problem with a server down indication. Autotune seemed to be running properly, so operators called an expert and he suggested Operations reboot a front end. This cleared the problem.

At 5:22 PM, CDF reported that they had turned off their silicon detector due to problems with the beam. Operations reported that this was related to a TeV RF station (TRF7) that had tripped off on Friday.

At 5:23 PM BRF6 tripped off. It reset.

Operations terminated store 5352 at 6:02 PM.

Operations turned on TRF7 per expert instructions at 6:12 PM.

Operations put the TeV into shot setup at 6:18 PM.

Operations reported at 8:19 PM that the Recycler had stashed four times during the last 12 hours. Accelerators status: the stack was at 42.0E10 and the stash was at 113.9E10.

Operations established store 5353 at 8:36 PM with an initial luminosity of 238.19E30.

Sunday April 15

The midnight shift began with Operations monitoring store 5353 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 4:33 AM, the Pelletron suffered a full discharge due to a vacuum burst. The Pelletron recovered and was back on by 5:30 AM.

At 5:56 AM, NuMI tripped off due to bad positions on the target. It reset.

At 7:41 AM, the Pelletron tripped off again. This time it would not reset. Operators talked with experts and then prepared the area for an access so they could reset the VME crate.

Operations reported at 8 AM that the Recycler had stashed four times during the last 12 hours.

At 8:29 AM, KRF3 tripped off due to a LCW flow indication. Operators reset the flow monitor.

Operations reported at 9:16 AM that with help from an expert they reset the VME and the expert restarted the front end, which got the Pelletron back up and running.

At 9:54 AM, Operations reported that the Booster LLRF room was warm. The Duty Electrician responded and reset the AC unit.

At 10:10 AM, LRF3 alarmed due to an RF vacuum problem. Operators resolved the problem.

At 10:31 AM, the Booster vacuum tripped off and caused a power supply and chopper to trip and vacuum valves to close. Operators managed to get everything back on about 25 minutes later.

At 11:04 AM, the I- Source vacuum gauge tripped off for the first of six times in five hours. It reset.

At 11:35 AM, BRF6 tripped off. It reset.

At 11:43 AM, LRF3 tripped off. It reset.

At 1:22 PM, KRF4 tripped off. The station reset, but the RF would not come up. Operations called an expert who thought the problem was due to the station’s charging power supply. With the help of the expert, operators reset the station. Beam resumed at 2:19 PM.

At 7:02 PM, TRF7 tripped off again and caused a small reduction in the store’s luminosity. Operations contacted an expert who said that the station’s power amplifier would have to be changed.

At 7:04 PM, NuMI tripped off. It reset.

Operations reported at 8 PM that the Recycler had stashed four times during the last 12 hours. Accelerator status: store 5353 was at 44.25E30, the stack was at 34.0E10, and the stash was at 383.2E10.

Operations terminated store 5353 at 9:05 PM.

TeV RF experts began their work on TRF7 at 9:06 PM.

At 11:02 PM, BRF6 tripped off. It reset.

At 11:18 PM, Operations reported that the Recycler had a damper problem that caused the beam emittance to blow up. Some antiprotons were lost.

TeV RF experts reported that they had completed their work on TRF7.

Operations put the TeV into a dry squeeze at 11:49 PM.

At 11:50 PM, the Duty Electrician reported that the AC unit for the Booster LLRF room had tripped off again.

Monday April 16

The midnight shift began with the next shot setup delayed due to Recycler damper problems and LLRF problems, with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:01 AM, Operations reported that the Recycler horizontal dampers were having serious problems.

At 1:13 AM, LRF3 tripped off. It reset.

At 1:29 AM, the Pbar lithium lens tripped off due to a power supply over temperature fault. Operators managed to get the lens back on by 2:12 AM.

At 2:27 AM, a NuMI power supply (V118) tripped off and would not reset. Operations called in an expert. The expert reported at 3:43 AM that he had reset the filter breaker and reported a possible door switch problem. NuMI beam resumed at 3:48 AM.

Operations reported at 4:09 AM that the Recycler emittances had come down.

Operations put the TeV into shot setup at 5:07 AM.

At 5:32 AM, LRF3 tripped off. It reset.

Operators made the Meson M01 area ready for access at 6:46 AM.

Operations established store 5355 at 7:11 AM with an initial luminosity of 230.35E30.

Operations reported at 8:09 AM that the Recycler had stashed twice during the last 12 hours.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow a MI study and a Pbar cooling alignment.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/16/2007    email Fermilab